OpenStudy (anonymous):

whats the y intercept of 7x+5y=7

hero (hero):

7x + 5y = 7 5y = 7 - 7x y = (7-7x)/5 y = 7/5 - 7x/5 y-int = 7/5 or 1.4

OpenStudy (anonymous):

thought so! thanks so much!

OpenStudy (anonymous):

solve the quadratic formula 2x^2-3x-7=0

hero (hero):

Use quad form

OpenStudy (anonymous):

The length of a rectangle is 3yrd more than twice its width, and the area of the rectangle is 77yrd^2 . Find the dimensions of the rectangle…